# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### 27 Feb 2024 v1

@@ 3 Convergence analysis of the PESS iterative process
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then $\Sigma^{-1}\mathcal{A}$ (normal-(((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
@@
Since the matrices $\Lambda_{1},\Lambda_{2},$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD, and further $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank. Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical. Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable. $\blacksquare$ By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix. Now, we establish an if and only if condition that precisely determines the convergence of the PESS iterative process.
@@
This implies $(2s-1)|\mu|^{2}+2\Re(\mu)^{2}>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive. $\blacksquare$ Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
@@
Again, we have $|\mu|<\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}>\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.9)
If we apply $s>0$, then we get the desired result from (3.9). $\blacksquare$ The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
@@ 4 Spectral distribution of the PESS preconditioned matrix
Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof. $\blacksquare$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, we have $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p},$ which can be written as
@@
  1. 1.
    Let $u=0$. Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD. Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector. Hence, $u\neq 0.$
  2. 2.
    Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
@@
  3. 3.
    The system of linear equations between (4.2a)-(4.2c) can be reformulated as
@@
In the following theorem, we provide sharper bounds for real eigenvalues of the PESS preconditioned matrix.
@@
$\lambda\in\left(0,\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}\right].$ (4.4)
@@
$\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)}\leq\lambda\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.7)
@@
$0<\lambda^{\pm}<\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.15)
@@
$\lambda>\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.17)
Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\lambda_{\max}(\Lambda_{1}^{-1}A).$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
@@
$\lambda\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{1+s\lambda_{\max}(\Lambda_{1}^{-1}A)}.$ (4.21)
Again, from (1) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.7) and (4.15), we obtained the desired bounds in (4.4) for $\lambda.$ Hence, the proof is concluded. $\blacksquare$ Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
@@
$\lambda\in\left(0,\frac{2\lambda_{\max}(A)}{\alpha+\lambda_{\max}(A)}\right].$ (4.22)
@@
Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.22) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$ In the next result, we discuss bounds for the real eigenvalues of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$.
@@
$\lambda\in\left(0,\frac{2\lambda_{\max}(P^{-1}A)}{\alpha+\lambda_{\max}(P^{-1}A)}\right].$ (4.23)
@@
Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.23) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$ The following result shows the bounds when $\lambda$ is a non-real eigenvalue.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds $(1).$   If $Cv=0,$ then $\lambda$ satisfies
$\frac{\lambda_{\min}(\Lambda^{-1}A)}{2+s\lambda_{\min}(\Lambda_{1}^{-1}A)}\leq|\lambda|\leq\sqrt{\frac{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)+s^{2}\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}}.$
$(2).$   If $Cv\neq 0,$ then real and imaginary part of $\lambda/(1-s\lambda)$ satisfies
$\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)\,\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{2\left(\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2},$
$\left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$
@@
$(1).$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds. We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$ Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular. Again, since $Cv=0,$ from (4.2c), and similar to (4.8)-(4.10), we obtain quadratic equation (4.11) in $\lambda$, which has complex solutions
@@
where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),$ we get the following bounds
$|\lambda^{\pm i}|^{2}$ $=$ $\Psi_{1}(p,q,s)\leq\Psi_{1}(\lambda_{\min}(\Lambda_{1}^{-1}A),\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),s)$ (4.26)
$=$ $\frac{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}{1+s\lambda_{\min}(\Lambda_{1}^{-1}A)+s^{2}\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})}.$
@@
$|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\lambda_{\min}(\Lambda_{1}^{-1}A))=\frac{\lambda_{\min}(\Lambda_{1}^{-1}A)}{2+s\lambda_{\min}(\Lambda_{1}^{-1}A)}.$ (4.27)
Hence, combining (4.26) and (4.27), proof for the part $(1).$ follows.
$(2).$   Next, consider the case $Cv\neq 0,$ then (4.2c) yields $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv,$ where $\mu=\lambda/(1-s\lambda).$ Substituting this in (4.2b), we get
@@
$|\Im(\mu)|\leq\sqrt{\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}.$ (4.35)
Furthermore, notice that $\Theta_{1}\geq\left(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A)\right)\Theta_{2},$ then from (4.33), we deduce
$-\Re(\mu)\geq\left(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A)\right)\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{\Re(\mu)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}.$ (4.36)
@@
$2\Re(\mu)\leq\lambda_{\max}(\Lambda_{1}^{-1}A)-\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\lambda_{\max}(\Lambda_{1}^{-1}A).$ (4.37)
@@
$2\Re(\mu)=\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\Re(\mu)\Theta_{2}-\Theta_{1})V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}\geq\lambda_{\min}(\Lambda_{1}^{-1}A)\,\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\mathcal{X}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v},$ (4.39)
@@
$-\theta_{i}<\Re(\mu)-\theta_{i}\leq\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\theta_{i},$
which further yields $(\Re(\mu)-\theta_{i})^{2}\leq\max\left\{\left(\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\theta_{i}\right)^{2},\theta_{i}^{2}\right\}.$ Hence, we get
$\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}$ $=$ $\displaystyle{\max_{i}}\{(\Re(\mu)-\lambda_{\min}(\Lambda_{1}^{-1}A))^{2},(\Re(\mu)-\lambda_{\max}(\Lambda_{1}^{-1}A))^{2}\}$ (4.40)
$\leq$ $\displaystyle{\max_{i}}\left\{\left(\frac{\lambda_{\max}(\Lambda_{1}^{-1}A)}{2}-\lambda_{\min}(\Lambda_{1}^{-1}A)\right)^{2},\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)\right\}$
$=$ $\lambda_{\max}^{2}(\Lambda_{1}^{-1}A).$
@@
$\mathcal{X}\geq\frac{I}{\lambda_{\max}^{2}(\Lambda_{1}^{-1}A)+\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})+\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C)}$ (4.41)

### 17 May 2024 v2

@@ 3 Convergence analysis of the PESS iterative process
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then $\Sigma^{-1}\mathcal{A}$ ((((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
@@
Since the matrices $\Lambda_{1},\Lambda_{2},$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD, and further $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank. Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical. Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable. $\blacksquare$
By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix. Now, we establish an if and only if a condition that precisely determines the convergence of the PESS iterative process.
@@
This implies $|\theta|<1$ if and only if $(2s-1)|\mu|^{2}+2\Re(\mu)>0.$ Since, $\vartheta(\mathcal{T})=\vartheta{\widetilde{(}\mathcal{T}}),$ the proof is conclusive. $\blacksquare$
Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
@@
Again, we have $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
If we apply $s>0$, then we get the desired result from (3.6). $\blacksquare$
The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
@@ 4 Spectral distribution of the PESS preconditioned matrix
Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof. $\blacksquare$
Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, we have $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p},$ which can be written as
@@ 3. $(3)$
$(1)$ Let $u=0$. Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD. Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector. Hence, $u\neq 0.$
$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
@@
$(3)$ The system of linear equations (4.2a)-(4.2c) can be reformulated as
@@
In the following theorem, we provide sharper bounds for real eigenvalues of the PESS preconditioned matrix. Before that, we introduce the following notation:
$\xi_{\max}:=\lambda_{\max}(\Lambda_{1}^{-1}A),\leavevmode\nobreak\ \xi_{\min}:=\lambda_{\min}(\Lambda_{1}^{-1}A),\leavevmode\nobreak\ \eta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T}),$ (4.4)
$\leavevmode\nobreak\ \eta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}B\Lambda_{1}^{-1}B^{T})\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \theta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}C^{T}\Lambda_{3}^{-1}C).$ (4.5)
@@
$\lambda\in\left(0,\frac{\xi_{\max}}{1+s\xi_{\max}}\right].$ (4.6)
@@
$\frac{\xi_{\min}}{1+s\xi_{\min}}\leq\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.9)
@@
$0<\lambda^{\pm}<\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.17)
@@
$\lambda>\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.19)
Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
@@
$\lambda\leq\frac{\xi_{\max}}{1+s\xi_{\max}}.$ (4.23)
Again, from ($(1)$) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.9) and (4.17), we obtained the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
Since the preconditioners ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{EGSS}$ are special cases of the PESS preconditioner, next, we obtain refined bounds for real eigenvalues of the SS and EGSS preconditioned matrices from Theorem 4.2.
@@
$\lambda\in\left(0,\frac{2\kappa_{\max}}{\alpha+\kappa_{\max}}\right],$ (4.24)
where $\kappa_{\max}=\lambda_{\max}(A).$
@@
Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$
In the next result, we discuss bounds for the real eigenvalues of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$.
@@
$\lambda\in\left(0,\frac{2\tilde{\kappa}_{\max}}{\alpha+\tilde{\kappa}_{\max}}\right],$ (4.25)
where $\tilde{\kappa}_{\max}=\lambda_{\max}(P^{-1}A).$
@@
Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$
The following result shows the bounds when $\lambda$ is a non-real eigenvalue.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds
  1. $(1)$
    If $Cv=0,$ then $\lambda$ satisfies
    $\frac{\xi_{\min}}{2+s\xi_{\min}}\leq|\lambda|\leq\sqrt{\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}}.$
  2. $(2)$
    If $Cv\neq 0,$ then real and imaginary part of $\lambda/(1-s\lambda)$ satisfies
    $\frac{\xi_{\min}\,\eta_{\min}}{2\left(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max}\right)}\leq\Re\left(\frac{\lambda}{1-s\lambda}\right)\leq\frac{\xi_{\max}}{2}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \left|\Im\left(\frac{\lambda}{1-s\lambda}\right)\right|\leq\sqrt{\eta_{\max}+\theta_{\max}},$
where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ is defined as in (4.4) and (4.5).
@@ 2. $(2)$
$(1)$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds. We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$
Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular. Again, since $Cv=0,$ from (4.2c), and similar to (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions
@@
where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds
$|\lambda^{\pm i}|^{2}=\Psi_{1}(p,q,s)$$\leq$ $\Psi_{1}(\xi_{\min},\eta_{\max},s)=\frac{\eta_{\max}}{1+s\xi_{\min}+s^{2}\eta_{\max}}.$ (4.28)
@@
$|\lambda^{\pm i}|=\Psi_{2}(p)\geq\Psi_{2}(\xi_{\min})=\frac{\xi_{\min}}{2+s\xi_{\min}}.$ (4.29)
Hence, combining (4.28) and (4.29), proof for the part $(1)$ follows. $(2)$   Next, consider the case $Cv\neq 0,$ then (4.2c) yields $w=\frac{1}{\mu}\Lambda_{3}^{-1}Cv,$ where $\mu=\lambda/(1-s\lambda).$ Substituting this in (4.2b), we get
@@
$|\Im(\mu)|\leq\sqrt{\eta_{\max}+\theta_{\max}}.$ (4.37)
Furthermore, notice that $\Theta_{1}\geq\left(\Re(\mu)-\xi_{\max}\right)\Theta_{2},$ then from (4.35), we deduce
$-\Re(\mu)\geq\left(\Re(\mu)-\xi_{\max}\right)\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\Theta_{2}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}+\frac{\Re(\mu)}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}.$ (4.38)
@@
$2\Re(\mu)\leq\xi_{\max}-\frac{\xi_{\max}}{|\mu|^{2}}\frac{v^{*}C^{T}\Lambda_{3}^{-1}Cv}{v^{*}\Lambda_{2}v}\leq\xi_{\max}.$ (4.39)
@@
$2\Re(\mu)=\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V(\Re(\mu)\Theta_{2}-\Theta_{1})V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v}\geq\xi_{\min}\,\frac{v^{*}B\Lambda_{1}^{-\frac{1}{2}}V\mathcal{X}V^{T}\Lambda_{1}^{-\frac{1}{2}}B^{T}v}{v^{*}\Lambda_{2}v},$ (4.41)
@@
$-\theta_{i}<\Re(\mu)-\theta_{i}\leq\frac{\xi_{\max}}{2}-\theta_{i},$
which further yields $(\Re(\mu)-\theta_{i})^{2}\leq\max\left\{\left(\frac{\xi_{\max}}{2}-\theta_{i}\right)^{2},\theta_{i}^{2}\right\}.$ Hence, we get
$\displaystyle{\max_{i}}(\Re(\mu)-\theta_{i})^{2}$ $=$ $\displaystyle{\max_{i}}\{(\Re(\mu)-\xi_{\min})^{2},(\Re(\mu)-\xi_{\max})^{2}\}$ (4.42)
$\leq$ $\displaystyle{\max_{i}}\left\{\left(\frac{\xi_{\max}}{2}-\xi_{\min}\right)^{2},\xi^{2}_{\max}\right\}=\xi^{2}_{\max}.$
@@
$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}$ (4.43)